hhanan:

Leonard makes $8.25 for every hour he works up to 40 hours a week. If he works overtime, he makes $14.88. Last week, he worked 40 hours plus n overtime hours and made a total of $456.75 on his check. How many overtime hours did he work? I think the answer is 9 but I am not sure

minustempo:

Write out an equation that represents how much money he makes 8.25*(40 hours) + 14.88*(n extra hours) = 456.75 move 8.25*40 to other side 14.88*n = 456.75-8.25*40 = 456.75-330 14.88n = 126.75

ILOVESPAGHETTI:

Alright so in this problem we would multiply $8.25 by 40 which equals $330 8.25 x 40 =330 So now we have 330, let's do a little method. A being our variable. A+330=456.75 So now just subtract 456.76 by 330 456.75-330.00 which equals 126.75 a=126.75

minustempo:

so n = 126.75/14.88 = 8.52 hours

ILOVESPAGHETTI:

@minustempo wrote:
so n = 126.75/14.88 = 8.52 hours
Correct!

hhanan:

@ilovespaghetti But it said round so then it would be 9?

ILOVESPAGHETTI:

@hhanan wrote:
@ilovespaghetti But it said round so then it would be 9?
well I didn't do all of the steps, it's 8.52 so you were really close. If it said round, then yes it's 9